# A Personalized Predictive Model that Jointly Optimizes Discrimination and Calibration

### v1

@@ 3 Simulation Studies
We conducted three simulation studies using R [21]. We first investigated the relationship between the size of the subpopulation, $M$, and the model discrimination measure, AUROC, to confirm the results of Lee et al. that as $M$ increases, the AUROC deteriorates [3]. We then investigated the relationship between calibration and the size of subpopulation, $M$. Finally, we applied our personalized predictive modelling algorithm to randomly generated data under a number of $\alpha$ values to determine how the value of $\alpha$ affects the optimal $M$ proportion found in the training/testing step of the algorithm.
@@ 3.1 Setup
We considered a sample size of $n=16000$. We simulated $20$ features, where $10$ were binary, $10$ were continuous, and the correlation between any two features was $r=0.2$. We simulated an outcome with a non-linear true outcome model, shown in Equation (5). This model was passed through an inverse logit function to obtain a probability, $\pi$, as shown in Equation (6). This probability was then used to simulate a Bernoulli outcome, $y_{k}$, where $Y_{k}\sim Bernoulli(\pi)$. The prevalence of the outcome was moderate, with values between 0.3 and 0.5 for all simulated datasets.
@@ 3.2 Relationship Between $M$ and the Performance Measures
We also tested three different weighting functions to investigate the effect of weighting the similar subpopulation on the performance of the PPMs. We considered uniform weights, (half-)tricube weights [22], and a purposely counter-intuitive weighting scheme which gives higher weights to individuals that are the least similar to the index patient. We found that the differences in model performance between the three weighting schemes were negligible, and concluded that the size of subpopulation has a much larger effect on the predictive performance of the PPM compared to the choice of weight function. This is a consistent finding with that in the curve estimation literature, where bandwidth size is much more important than neighbourhood weighting.
@@ 4 Data Analysis
We applied our proposed algorithm to a dataset from the eICU Collaborative Research Database [15] to predict mortality of patients with sepsis. The eICU database is a large multi-center critical care database from the United States and is sourced from the eICU Telehealth Program, which is a model of care where remote providers monitor patients continuously. The database consists of 200,859 unit encounters for 139,367 unique patients. These patients were admitted between 2014 and 2015 at 208 hospitals located throughout the USA. The database is deidentified, and includes tables with vital sign measurements, nurse’s notes, severity of illness scoring results, diagnoses, treatment information, care plan documentation, and more. One specific table in the eICU database, vitalPeriodic, includes data which is interfaced from a bedside monitor at one-minute averages, but is archived in the vitalPeriodic table as five-minute median values. Data is publicly available via a PhysioNet repository. However, to access, one must complete a training course in research with human subjects.
@@
Summary statistics are often used to summarize longitudinal data in analyses of ICU data, such as the vital signs in the case of this data analysis. Such summary statistics might include the mean or median of the repeated measures for a given vital sign, as well as a minimum and maximum value. However, just including summary statistics does not consider the temporal aspect of vital signs. Thus, due to the dense curve-like nature of the vital signs, we use functional principal components (FPCs) to identify each of their dominant modes of variation, specifically for the vital signs in the first 24-hours of a patient’s stay in the ICU. We applied FPCA to the vital sign data using the Principal Analysis by Conditional Estimation (PACE) algorithm provided in the R package fdapace [23]. The PACE algorithm uses eigenfunctions to expand the trajectory of a longitudinal feature for a given patient, as shown in Equation 7. The first $r$ eigenfunctions ($\hat{\phi}_{q}(t),q=1,...,r)$ are used in this expansion, where $\hat{\mu}(t)$ is the estimated mean function of the feature and $\hat{\xi}_{kq}$ represent the FPCs. We extracted the first three FPCs for each of the vital signs, which cumulatively explained at least 95% of the total variation in the trajectory of each given vital sign; a different dataset for these vital signs, or even different vital signs in this dataset, may have required a different number of FPCs to account for at least 95% variation.. Our resulting dataset thus consisted of 17 features, including the first three FPC scores for each of the three vital signs.
@@
When testing the algorithm, we considered $\alpha=\{0.475,0.49,0.5,0.51,0.55,0.6,0.750,0.9,0.95\}$. We conducted 20 repeated 10-fold CV in the training/testing step of the algorithm. Similar to the simulation study, our outcome in this case is binary, and we fit a logistic regression model to the data. We weighted the subpopulation uniformly. The prevalence of outcome in this dataset is about 0.107, which is much lower than that of our simulation study (about 0.4). Because of this, we calculated the area under the precision recall curve (AUPRC) as a measure of discrimination in addition to the AUROC, as the latter can be misleading when the prevalence of outcome is highly skewed [24]. Specifically, in these cases, the AUROC tends to be too optimistic.
@@ 5 Discussion
While there have been algorithms proposed to fit personalized predictive models (PPMs), there is lack of emphasis in the literature on models that are well-calibrated. We thus proposed an algorithm which fits a PPM using the optimal size of subpopulation that considers both discrimination and calibration. This is done through a proposed loss function that allows for flexibility in emphasizing one performance measure over another, as desired.
@@
The trade-off between discrimination and calibration when choosing an $M$ value was the motivation behind our proposed algorithm which fits a PPM using an $M$ that considers both performance measures. Our proposed loss function, an extension of the Brier Score, uses the mixture term $\alpha$ to emphasize either performance measure. When the mixture term, $\alpha$, is set to 0.5, then the proposed mixture loss function is proportional to Brier Score. Thus, they attain their minimum at the same value of $M$. Through simulation studies, we found that as we increased $\alpha$, the optimal $M$ proportion found in the training/testing step of the algorithm also increased. In other words, as we put more emphasis on the calibration term in our proposed mixture loss function, the optimal proportion of subpopulation to fit the PPM also increased. However, this rate of increase was not consistent. The largest rate of change we see is between $\alpha=0.75$ and $\alpha=0.85$. Before this, the changes in the optimal $M$ proportion between subsequent $\alpha$ values in very gradual. These results were robust across the five randomly generated datasets. In our study, the best performing model was when $\alpha=0.58$, which means slightly more emphasis was put on calibration compared to discrimination in the mixture loss function. The main purpose of this study was to show that researchers can use the Brier Score as the loss function, but the resulting model will not necessarily be optimal in a general setting. What is optimal will depend on the user and what they care about, whether that is a model that has better discrimination over calibration or vice versa (or a balance between the two), and our next step is to suggest criterion in choosing this $\alpha$ value so the user can can obtain the model performance desired. We were limited in the number of $\alpha$ values below 0.5 we could test, as these $\alpha$ values resulted in very small optimal $M$ proportions. When we applied the optimal $M$ proportion to the validation step, the subpopulation used for training the PPM was not a sufficient size.
@@
Our study has a few limitations. Our algorithm is very computationally extensive. We employed the Rcpp package [25], along with parallelization, which sped up the running time significantly, but there is still much room for improvement. If the algorithm can be adjusted to become computationally efficient, other statistical models such as random forest or gradient boosting machines can be tested to see if our results translate to other models. Further, in both our simulation study and the data analysis, we considered a relatively small number of features. When working with EHR data, there may be hundreds of features available when fitting prediction models. Thus, a natural next step is to include a variable selection step to our proposed algorithm. Finally, we do not know if the relationships we have shown are specific to the data we considered. Eventually, we want to provide researchers with criterion on choosing an $\alpha$ value, but more work needs to be put into investigating various data scenarios before we can provide more informed advice.

### v2

@@ 3 Simulation Studies
We conducted three simulation studies using R [R Core Team, 2020]. We first investigated the relationship between the size of the subpopulation, $M$, and the model discrimination measure, AUROC, to confirm the results of Lee et al. that as $M$ increases, the AUROC deteriorates. [Lee et al., 2015] We then investigated the relationship between calibration and the size of subpopulation, $M$. Finally, we applied our personalized predictive modelling algorithm to randomly generated data under a number of $\alpha$ values to determine how the value of $\alpha$ affects the optimal $M$ proportion found in the training/testing step of the algorithm.
@@ 3.1 Setup
We considered a sample size of $n=16000$. We simulated $20$ features, where $10$ were binary, $10$ were continuous, and the correlation between any two features was $r=0.2$. We simulated an outcome with a non-linear true outcome model, shown in Equation (6). This model was passed through an inverse logit function to obtain a probability, $\pi$, as shown in Equation (7). This probability was then used to simulate a Bernoulli outcome, $y_{k}$, where $Y_{k}\sim Bernoulli(\pi)$. Note that the simulated outcome of each index patient in the testing dataset, $y_{I}$, was generated in the same way. The prevalence of the outcome was moderate, with values between 0.3 and 0.5 for all simulated datasets.
@@
In practice, this outcome model is usually unknown, thus in our simulation study, we do not fit the model described in Equation (6) to the index patient to obtain a prediction. Instead, we fit a logistic regression model which considers all $20$ simulated features, but no interactions.
@@ 3.2 Relationship Between $M$ and the Performance Measures
We also tested three different weighting functions to investigate the effect of weighting the similar subpopulation on the performance of the PPMs. We considered uniform weights, (half-)tricube weights [Cleveland, 1979], and a purposely counter-intuitive weighting scheme which gives higher weights to individuals that are the least similar to the index patient. We found that the differences in model performance between the three weighting schemes were negligible, and concluded that the size of subpopulation has a much larger effect on the predictive performance of the PPM compared to the choice of weight function. This is a consistent finding with that in the curve estimation literature, where bandwidth size is much more important than neighbourhood weighting.
@@ 4 Data Analysis
We applied our proposed algorithm to a dataset from the eICU Collaborative Research Database [Pollard et al., 2018] to predict mortality of patients with sepsis. The eICU database is a large multi-center critical care database from the United States and is sourced from the eICU Telehealth Program, which is a model of care where remote providers monitor patients continuously. The database consists of 200,859 unit encounters for 139,367 unique patients. These patients were admitted between 2014 and 2015 at 208 hospitals located throughout the USA. The database is deidentified, and includes tables with vital sign measurements, nurse’s notes, severity of illness scoring results, diagnoses, treatment information, care plan documentation, and more. One specific table in the eICU database, vitalPeriodic, includes data which is interfaced from a bedside monitor at one-minute averages, but is archived in the vitalPeriodic table as five-minute median values. Data is publicly available via a PhysioNet repository. However, to access, one must complete a training course in research with human subjects.
@@
Summary statistics are often used to summarize longitudinal data in analyses of ICU data, such as the vital signs in the case of this data analysis. Such summary statistics might include the mean or median of the repeated measures for a given vital sign, as well as a minimum and maximum value. However, just including summary statistics does not consider the temporal aspect of vital signs. Thus, due to the dense curve-like nature of the vital signs, we use functional principal components (FPCs) to identify each of their dominant modes of variation, specifically for the vital signs in the first 24-hours of a patient’s stay in the ICU. We applied FPCA to the vital sign data using the Principal Analysis by Conditional Estimation (PACE) algorithm provided in the R package fdapace. [Carroll et al., 2020] The PACE algorithm uses eigenfunctions to expand the trajectory of a longitudinal feature for a given patient, as shown in Equation 8. The first $r$ eigenfunctions ($\hat{\phi}_{q}(t),q=1,...,r)$ are used in this expansion, where $\hat{\mu}(t)$ is the estimated mean function of the feature and $\hat{\xi}_{kq}$ represent the FPCs. We extracted the first three FPCs for each of the vital signs, which cumulatively explained at least 95% of the total variation in the trajectory of each given vital sign; a different dataset for these vital signs, or even different vital signs in this dataset, may have required a different number of FPCs to account for at least 95% variation.. Our resulting dataset thus consisted of 13 features, including the first three FPC scores for each of the three vital signs.
@@
When testing the algorithm, we considered $\alpha=\{0.475,0.49,0.5,0.51,0.55,0.6,0.750,0.9,0.95\}$. We conducted 20 repeated 10-fold CV in the training/testing step of the algorithm. Similar to the simulation study, our outcome in this case is binary, and we fit a logistic regression model to the data. We weighted the subpopulation uniformly. The prevalence of outcome in this dataset is about 0.107, which is much lower than that of our simulation study (about 0.4). Because of this, we calculated the area under the precision recall curve (AUPRC) as a measure of discrimination in addition to the AUROC, as the latter can be misleading when the prevalence of outcome is highly skewed. [Davis and Goadrich, 2006] Specifically, in these cases, the AUROC tends to be too optimistic.
@@ 5 Discussion
While there have been algorithms proposed to fit personalized predictive models (PPMs), there is lack of emphasis in the literature on models that are well-calibrated. We thus proposed an algorithm which fits a PPM using the optimal size of subpopulation that considers both discrimination and calibration. This is done through a novel loss function that we propose which allows for flexibility in emphasizing one performance measure over another, as desired.
@@
The trade-off between discrimination and calibration when choosing an $M$ value was the motivation behind our proposed algorithm which fits a PPM using an $M$ that considers both performance measures. Our proposed loss function, an extension of the Brier Score, uses the mixture term $\alpha$ to emphasize either performance measure. When the mixture term, $\alpha$, is set to 0.5, then the proposed mixture loss function is proportional to Brier Score. Thus, they attain their minimum at the same value of $M$. Through simulation studies, we found that as we increased $\alpha$, the optimal $M$ proportion found in the training/testing step of the algorithm also increased. In other words, as we put more emphasis on the calibration term in our proposed mixture loss function, the optimal proportion of subpopulation to fit the PPM also increased. However, this rate of increase was not consistent. The largest rate of change we see is between $\alpha=0.75$ and $\alpha=0.85$. Before this, the changes in the optimal $M$ proportion between subsequent $\alpha$ values in very gradual. These results were robust across the five randomly generated datasets. In our study, the best performing model was when $\alpha=0.58$, which means slightly more emphasis was put on calibration compared to discrimination in the mixture loss function. The main purpose of this study was to show that researchers can use the Brier Score as the loss function, but the resulting model will not necessarily be optimal in a general setting. What is optimal will depend on the user and what they care about, whether that is a model that has better discrimination over calibration or vice versa (or a balance between the two), and our next step is to suggest criterion in choosing this $\alpha$ value so the user can can obtain the model performance desired. The $\alpha$ value can also be treated as a tuning parameter in our proposed algorithm, however, this would increase the already intensive computational burden. We were limited in the number of $\alpha$ values below 0.5 we could test, as these $\alpha$ values resulted in very small optimal $M$ proportions. When we applied the optimal $M$ proportion to the validation step, the subpopulation used for training the PPM was not a sufficient size.
@@
Our study has a few limitations. First, we acknowledge that implementing this procedure for a new index patient will require that the index patient comes from (roughly) the same population of patients on which the $M$ was originally tuned. This may need to be judged on a case-by-case basis. Certainly, we would not want to use an $M$ derived from a different country’s set of hospitals at a different point in time, thus some of the same considerations for externally validating existing prediction models would need to be considered here. Further, our algorithm is very computationally extensive. We employed the Rcpp package [Eddelbuettel and François, 2011], along with parallelization, which sped up the running time significantly, but there is still much room for improvement. If the algorithm can be adjusted to become computationally efficient, other statistical models such as random forest or gradient boosting machines can be tested to see if our results translate to other models, as changing the model that is fit could have an effect on the chosen $M$. Specifically, we wish to investigate how much of an effect the complexity and flexibility of the models fit has on the chosen $M$ value. In addition, in the simulation study and the data analysis we conduct a complete case analysis for simplicity to demonstrate the use of our proposed algorithm, but in practice it is recommended to apply imputation methods to handle the missing data. Implementing such a method, such as multiple imputation, will add another layer of complexity to the algorithm, and how to minimize this added complexity is an area of future work. We also note that the missingness could be be informative regarding similarity between patients, and this idea should also be explored. For example, patients with consistent missing data patterns across different covariates of interest, such as lab tests, could be deemed similar. Another limitation to the studies conducted in this paper is that we considered a relatively small number of features. When working with EHR data, there may be hundreds of features available when fitting prediction models. Thus, a natural next step is to include a variable selection step to our proposed algorithm. Finally, we do not know if the relationships between $M$ and calibration, as well as $M$ and discrimination, that we have shown in this study are specific to the data we considered. Eventually, we want to provide researchers with criterion on choosing an $\alpha$ value, but more work needs to be put into investigating various data scenarios before we can provide more informed advice.